OpenStudy (anonymous):

What is the range of the graph of y = 3(x – 3)^ 2 + 4?

jimthompson5910 (jim_thompson5910):

The range of \[\large y=a(x-h)^2+k\] is \[(-\infty,k]\] if \[a<0\] ------------- The range of \[\large y=a(x-h)^2+k\] is \[[k,\infty)\] if \[a>0\]

OpenStudy (anonymous):

yeah i have no idea what this is, can you just give me the answer so i can finish this exam

jimthompson5910 (jim_thompson5910):

That's the general formula for a quadratic in vertex form, which is what you have there. In this case, h=3 and k=4 What is the value of 'a'?

OpenStudy (anonymous):

ooooh, i get it never mind, thx alot, this was really helpful